A jeweler is commissioned to create a piece that will weigh 14oz. And consist of 75% gold. Two alloys are available that are 60% and 80% gold respectively. How much of each alloy should be used?
Answer by lwsshak3(11628)

A jeweler is commissioned to create a piece that will weigh 14oz. And consist of 75% gold. Two alloys are available that are 60% and 80% gold respectively. How much of each alloy should be used?
let x= amt of 60% alloy used
14-x= amt of 80% alloy used
60%x+80%(14-x)=75%*14
.60x+11.2-.80x=10.5
.2x=.70
x=3.5
14-x=10.5
amt of60% alloy used=3.5 oz
amt of 80% alloy used=10.5 oz